OpenStudy (anonymous):

Please check me answer! Which is true about the solutions to x2 + 2x−8=0? a. No real solutions b. Two identical rational solutions c. Two different rational solutions d. Two irrational solutions Is the answer C correct?

OpenStudy (anonymous):

To determine the types and number of roots of a quadratic equation, examine the discriminant: • If b² - 4ac > 0 then there are two real (unique) roots • If b² - 4ac = 0 then there is one (double) real root • If b² - 4ac < 0 then there are two imaginary roots In this case: x^2 + 2x - 8 = 0 • (2)² - 4(1)(-8) = 36 > 0 thus there are two real (unique) roots And since √36 = 6, both roots are rational.

OpenStudy (kira_yamato):

x^2 + 2x−8=0 \[x = \frac{-2 ± \sqrt{2^2 - 4(1)(-8)}}{2(1)} = \frac{-2 ± \sqrt{36}}{2}=-1 ± 6\] x = -1 + 6 = 5 or x = -1 - 6 = -7

OpenStudy (kira_yamato):

Therefore the answer is c. Two different rational solutions

OpenStudy (anonymous):

Thank you guys!